ROTATIONAL ANALYSIS OF LASER INDUCED FLUORESCENCE DATA FOR Rh^{16}O and Rh^{18}O

\underline{R. F. HEUFF} and W. J. BALFOUR, Department of Chemistry, University of Victoria, Victoria, BC Canada, V8W 3P6.

We recently reported the laser induced fluorescence^{a} and dispersed fluorescence^{b} spectra for rhodium monoxide. Molecular constants were derived from our rotational analysis of the four 0-0 bands of two ^{2}Π ← X^{4}Σ transitions using jet-cooled (J = 15) high-resolution data. We have now extended the analysis to include data from 16 Rh^{16}O medium-resolution sub-bands (J = 35) arising from the v = 0 level of the ground state. In addition, emission plates from Stockholm, recorded in the 1960s have been examined, extending the analysis to much higher J. Several bands for Rh^{18}O have also been investigated. Spectroscopic parameters from some 3000 assigned lines will be discussed.

aR. F. Heuff, W. J. Balfour and A. G. Adam, J. Mol. Spec. 216 136 (2002).

bR. H. Jensen, S. G. Fougere and W. J. Balfour, Chem. Phys. Lett. 370 106 (2003).